# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis

### v3

@@ 5.2.2 Operator Ratios From Proper Positive Semidefinite Functions
We extend the concept of operator ratios (defined at Eq. (5)) for which the arguments are positive semidefinite functions rather than positive definite functions. The arguments can also be functions which take infinite values. So, let us denote by $\mathcal{PSD}$ the set of nonzero proper positive semidefinite functions from $\mathbb{R}^{d}$ to $[0,+\infty]$ i.e. :
@@
We are able to extend the operator ratio function to $\mathcal{PSD}$ functions.
@@
Let $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ and $k\in\mathbb{N}$. We associate to $U$ and $k$ the operator ratio function $\overline{N_{U^{k}}}:\mathcal{PSD}\mapsto[0,+\infty]$ defined as follows:
$\overline{N_{U^{k}}}(P):=\sup_{P(x)\in\mathbb{R}_{+}^{*}}\dfrac{P(U^{k}(x))}{P(x)}$
As for $N_{U^{k}}$, we will simply write $\overline{N_{U}}$ for $k=1$.
Recall that $\mathrm{Id}$ denotes the identity function $\mathbb{R}^{d}\ni x:\mapsto x$. It is straightforward to see that, if $P\in\mathcal{PSD}$, $\overline{N_{\mathrm{Id}}}(P)=1$.
Example 6.
We proved at Example 5 that:
$\overline{N_{T}}(V)=\sup_{\{x\in\mathbb{R}^{2}:V(x)>0\}}\dfrac{V(T(x))}{V(x)}=\max_{x\in\operatorname{Vect}\{(2,1)^{\intercal}\}\backslash\{0\}}\dfrac{V(T(x))}{V(x)}=4/9.$
The operator function is well-defined but the inequality $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)^{k}$ can fail to be true as we can see at Example 7.
@@
Let us consider the function $U:x\mapsto-x$ and $P:x\mapsto\max\{x,0\}$. We have $\overline{N_{U}}(P)=\sup_{P(x)>0}P(U(x))/P(x)=\sup_{x>0}P(-x)/x=\sup_{x>0}0/x=0$ whereas $\overline{N_{U^{2}}}(P)=\sup_{x>0}P(x)/x=\sup_{x>0}x/x=1$. Finally, we have $\overline{N_{U^{2}}}(P)>\overline{N_{U}}(P)^{2}$.
@@
$\mathcal{D}\left(U\right):=\{P\in\mathcal{PSD}:\forall\,x\in\mathbb{R}^{d},\ P(U(x))=0\implies P(U^{2}(x))=0\}$ (8)
@@
Let consider $U(x)=-x$ and $P(x)=\max\{x,0\}$. Let $x$ be a strictly positive real. Then, $P(U(x))=P(-x)=0$. However, $P(U^{2}(x))=P(x)=x>0$. So $P\notin\mathcal{D}\left(U\right)$.
@@
For $k=0$, the result holds as $\overline{N_{\mathrm{Id}}}(P)=1$ if there exists $x\in\mathbb{R}^{d}$ such that $P(x)>0$ and 0 otherwise. If $\overline{N_{U}}(P)$ is finite $(\overline{N_{U}}(P))^{0}$ is defined and equal to 1.
Now let $k\in\mathbb{N}^{*}$. If $\overline{N_{U}}(P)=+\infty$, the inequality is obviously true. Suppose now that $\overline{N_{U}}(P)<+\infty$. From the defintion of $\overline{N_{U}}(P)$, we have for all $x\in\mathbb{R}^{d}$ and all $j\in\mathbb{N}^{*}$ such that $P(U^{j-1}(x))\in\mathbb{R}_{+}^{*}$, $P(U^{j}(x))\leq\overline{N_{U}}(P)P(U^{j-1}(x))$. From the fact that $P\in\mathcal{D}\left(U\right)$, we also have for all $x\in\mathbb{R}^{d}$ and all $j\in\mathbb{N}^{*}$ such that $P(U^{j}(x))=0$, $P(U^{j+1}(x))\leq\overline{N_{U}}(P)P(U^{j}(x))$ as it implies that $P(U^{j+1}(x))=0$. Finally, we have, for all $x\in\mathbb{R}^{d}$ such that $P(x)>0$, $P(U^{k}(x))\leq\overline{N_{U}}(P)P(U^{k-1}(x))$. By dividing by $P(x)$ and taking the supremum, we obtain $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)\overline{N_{U^{k-1}}}(P)$ and we conclude by recurrence. ∎
@@
First $V^{-1}(\mathbb{R}_{+}^{*})$ is nonempty from the second statement of Def. 8. The third statement implies that $V(T^{2}(x))\leq\lambda V(T(x))$ for all $x\in\mathbb{R}^{d}$ for some $\lambda\in(0,1)$ and thus $V(T^{2}(x))>0$ implies that $V(T(x))>0$. We conclude that $V\in\mathcal{D}\left(T\right)$. The third statement also implies that $V(T(x))\leq\lambda V(x)$ for all $x\in\operatorname{dom}(V)$ for some $\lambda\in(0,1)$ and thus $N_{T}(V)\leq\lambda<1$. It can be equal to 0 if and only if for all $x$ such that $V(x)>0$, $V(T(x))=0$. ∎
To prove that a function $V\in\mathcal{PSD}$ is an Opt-Lyapunov for $X^{\mathrm{in}},T)$, we could use any value greater $\overline{N_{T}}(V)$. The computation of $\overline{N_{T}}(V)$ ensures to take the best value possible. It is important in the context of using the function $\mathfrak{F}_{u}(k,h,\beta)$ with an Opt-Lyapunov function since $\beta\mapsto\mathfrak{F}_{u}(k,h,\beta)$ is decreasing (Prop. 12).
@@
If there exists a Lyapunov function for $U$, then $U(0)=0$. From the double inequalities $\alpha\circ\|\cdot\|\leq V\leq\beta\circ\|\cdot\|$ and thus $V(x)=0$ iff $x=0$. Thus if $V(U^{2}(x))>0$ then $U^{2}(x)\neq 0$ and thus $U(x)\neq 0$ hence $V(U(x))>0$ and $V\in\mathcal{D}\left(U\right)$. The equality between $\overline{N_{U}}(V)$ and $N_{U}(V)$ follows readily from the fact that $\{y\in\mathbb{R}^{d}:V(y)>0\}=\mathbb{R}^{d}\backslash\{0\}$. ∎
5.2.3 $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility and $(X^{\mathrm{in}},T,\varphi)$-compatibility
In the proof of Th. 7, the construction of $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ from a classical Lyapunov function $V$ is based on the existence of a function $\alpha\in\mathcal{K}_{\infty}$ such that $\varphi\leq\alpha\circ V$. The existence of this latter comes from the fact that in Th. 7, $\varphi$ is continuous and, by definition of $V$, $\alpha^{\prime}\circ\|\cdot\|\leq V$ for some $\alpha^{\prime}\in\mathcal{K}_{\infty}$. However, when $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, $V$ is not necessarily positive definite and we cannot guarantee the existence of $\alpha\in\mathcal{K}_{\infty}$ such that $\alpha\circ\varphi\leq V$. We have to use a different class of function to find a similar relation between $\varphi$ and $V$. First, we propose to restrict the interval where a function $\alpha$ should be continuous and strictly increasing. Indeed, the analysis only concerns the values $\{\varphi(T^{k}(x)),k\in\mathbb{N},x\in X^{\mathrm{in}}\}$ and $\{V(x),x\in X^{\mathrm{in}}\}$. Hence, we introduce the following interval :
@@
    For all $k\in\mathbb{N}$, $\nu_{k}\in I_{X^{\mathrm{in}},T}^{\varphi}$. The values, if they are finite, $\limsup_{n\in\mathbb{N}}\nu_{n}$ and $\liminf_{n\in\mathbb{N}}\nu_{n}$ also belong to $I_{X^{\mathrm{in}},T}^{\varphi}$.
@@
Let $k\in\mathbb{N}$. There exists for all $n\in\mathbb{N}$, $x_{n}\in X^{\mathrm{in}}$ such that $\nu_{k}-1/n\leq\varphi(T^{k}(x_{n}))\leq\nu_{k}$ and thus the sequence $(\varphi(T^{k}(x_{n})))_{n\in\mathbb{N}}$ converges to $\nu_{k}$ and thus $\nu_{k}\in I_{X^{\mathrm{in}},T}^{\varphi}$. The values, if they are finite, $\limsup_{k\in\mathbb{N}}\nu_{k}$ and $\liminf_{k\in\mathbb{N}}\nu_{k}$ are limits of subsequences of $(\nu_{k})_{k\in\mathbb{N}}$ and they belongs to $I_{X^{\mathrm{in}},T}^{\varphi}$. From Prop. 1, if $\limsup_{k\in\mathbb{N}}\nu_{k}$ (resp. $\liminf_{k\in\mathbb{N}}\nu_{k}$) is infinite, the value $\sup_{k\in\mathbb{N}}\nu_{k}$ (resp. $\inf_{k\in\mathbb{N}}\nu_{k}$ is equal to +$\infty$+ ∞ (resp. $\infty$- ∞).
Now, if one of the value $\inf_{k\in\mathbb{N}}\nu_{k}$ or $\sup_{k\in\mathbb{N}}\nu_{k}$ is not finite, the infinity passes to the related bound of $I_{X^{\mathrm{in}},T}^{\varphi}$. Now if those values are both finite, we can extract sequences in $I_{X^{\mathrm{in}},T}^{\varphi}$ which converge respectively to $\inf_{m\in\mathbb{N}}\nu_{m}$ and $\sup_{m\in\mathbb{N}}\nu_{m}$. For example, we can define $\varphi(T^{k_{n}}(x_{k_{n}}))\subset I_{X^{\mathrm{in}},T}^{\varphi}$ such that for all $n\in\mathbb{N}$, $\inf_{m\in\mathbb{N}}\nu_{m}-1/n\leq\nu_{k_{n}}-1/n\leq\varphi(T^{k_{n}}(x_{k_{n}}))\leq\nu_{k_{n}}\leq\inf_{m\in\mathbb{N}}\nu_{m}+1/n$. This sequence converge to $\inf_{m\in\mathbb{N}}\nu_{m}$. ∎
Let $V$ be an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. To pass from $\alpha\circ\varphi\leq V$ to $\varphi\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\alpha\circ V$, $\alpha$ has to be surjective and strictly increasing as a function of $J\to\{V(x),x\in X^{\mathrm{in}}\}\subseteq[0,1]$. From the earlier discussion, $J$ must also contain $I_{X^{\mathrm{in}},T}^{\varphi}$. Therefore, we define the following set of functions:
@@
Let us consider $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and $I$ such that $\alpha(I)=[0,1]$. As $\alpha$ is continuous and strictly increasing on $I$, ${{\alpha}_{\left(I\right)}^{-1}}$ exists and is also continuous and strictly increasing from $[0,1]\to I$. Hence, $I$ is compact. Moreover, such interval $I$ is unique. Otherwise, consider $I,J\in\mathrm{SIC}(\alpha)$ for which we have $I\neq J$ and $\alpha(I)=\alpha(J)=[0,1]$. Suppose there exists $x\in I$ such that $x\leq\min J$. We have $\alpha(x)<\alpha(\min J)=0$ but $x\in I$ and $\alpha(x)\geq\alpha(\min I)=0$. ∎
We present some useful properties of functions belonging to $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and their relation with $\nu$.
@@
For all $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that there exists an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, $V$, satisfying $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, we have $\alpha(\limsup_{k\to+\infty}\nu_{k})\leq 0$.
@@
Let $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. By assumption, $\alpha(\varphi(T^{k}(x)))\leq V(\varphi(T^{k}(x)))\leq\lambda^{k}V(x)$. Taking the sup over $X^{\mathrm{in}}$ and from the increasing condition and the continuity of $\alpha$, $\alpha(\nu_{k})\leq\lambda^{k}\overline{V_{X^{\mathrm{in}}}}$. Taking the limsup, we get $\limsup_{k\mapsto+\infty}\alpha(\nu_{k})=\alpha(\limsup_{k\mapsto+\infty}\nu_{k})\leq 0$. ∎
@@
Suppose there exist $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ $V$ such that $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. Then:
@@
The first statement follows readily from Prop. 22. The second statement follows from Prop. 3 and Prop. 22, since we have, if ${\mathrm{R}}_{\nu}^{s}=\emptyset$, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\sup_{k\in\mathbb{N}}\nu_{k}=\limsup_{n\to+\infty}\nu_{n}$. ∎
The set $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ defines the class of functions that we must consider to obtain relations $\alpha\circ\varphi\leq V$ when $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. Another important consideration to take into account is usefulness of functions $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ based on $\alpha$. An Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is always nonnegative and if $\alpha(\nu_{k})\leq 0$ for all $k\in\mathbb{N}$, the inequality is obvious. Then $\alpha(\nu_{k})$ must be strictly positive for some $k\in\mathbb{N}$. By Prop. 22 and Corollary 5, we have automatically $k\in{\mathrm{R}}_{\nu}^{s}$. We propose a special definition for those particular functions $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$.

### v4

@@ 5.1.2 Operator Ratios From Proper Positive Semidefinite Functions
We extend the concept of operator ratios developed in [Adj25a] (specialized at Eq. (17) for Lyapunov functions) for which the arguments are positive semidefinite functions rather than positive definite functions. The arguments can also be functions that take infinite values. Therefore, let us denote by $\mathcal{PSD}$ the set of nonzero proper positive semidefinite functions from $\mathbb{R}^{d}$ to $0,+\infty]$, that is,
@@
We can extend the operator ratio function to $\mathcal{PSD}$ functions.
@@
Let $U:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$. We associate to $U$ the operator ratio function $\overline{N_{U}}:\mathcal{PSD}\mapsto[0,+\infty]$ defined as follows:
$\mathcal{PSD}\ni P\mapsto\overline{N_{U}}(P):=\sup_{P(x)\in\mathbb{R}_{+}^{*}}\dfrac{P(U(x))}{P(x)}$
Recall that $\mathrm{Id}$ denotes the identity function $\mathbb{R}^{d}\ni x:\mapsto x$. It is straightforward to see that if $P\in\mathcal{PSD}$, $\overline{N_{\mathrm{Id}}}(P)=1$.
The operator function is well-defined, but the inequality $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)^{k}$ can fail to be true, as shown in Example 5.
@@
Let us consider the functions $U:x\mapsto-x$ and $P:x\mapsto\max\{x,0\}$. We have $\overline{N_{U}}(P)=\sup_{P(x)>0}P(U(x))/P(x)=\sup_{x>0}P(-x)/x=\sup_{x>0}0/x=0$ whereas $\overline{N_{U^{2}}}(P)=\sup_{x>0}P(x)/x=\sup_{x>0}x/x=1$. Finally, we have $\overline{N_{U^{2}}}(P)>\overline{N_{U}}(P)^{2}$.
@@
$\mathcal{D}\left(U\right):=\{P\in\mathcal{PSD}:P(U(x))=0\implies P(U^{2}(x))=0\}$ (34)
@@
The functions $U$ and $P$ are the ones defined in Example 5. Let $x>0$ be strictly positive. Then, P(U(x))=P(x)=0P(U(x))=P(-x)=0. However, $P(U^{2}(x))=P(x)=x>0$. So $P\notin\mathcal{D}\left(U\right)$.
@@
For $k=0$, the result holds as $\overline{N_{\mathrm{Id}}}(P)=1$ and there exists $x\in\mathbb{R}^{d}$ such that $P(x)>0$. If $\overline{N_{U}}(P)$ is finite, $(\overline{N_{U}}(P))^{0}$ is defined and equal to 1.
Now let $k\in\mathbb{N}^{*}$. If $\overline{N_{U}}(P)=+\infty$, then the inequality is obviously true. Suppose now that $\overline{N_{U}}(P)<+\infty$. From the defintion of $\overline{N_{U}}(P)$, we have for all $x\in\mathbb{R}^{d}$ and all $j\in\mathbb{N}^{*}$ such that $P(U^{j-1}(x))\in\mathbb{R}_{+}^{*}$, $P(U^{j}(x))\leq\overline{N_{U}}(P)P(U^{j-1}(x))$. From the fact that $P\in\mathcal{D}\left(U\right)$, we also have for all $x\in\mathbb{R}^{d}$ and all $j\in\mathbb{N}^{*}$ such that $P(U^{j}(x))=0$, $P(U^{j+1}(x))\leq\overline{N_{U}}(P)P(U^{j}(x))$ as it implies that $P(U^{j+1}(x))=0$. Finally, we have, for all $x\in\mathbb{R}^{d}$ such that $P(x)>0$, $P(U^{k}(x))\leq\overline{N_{U}}(P)P(U^{k-1}(x))$. By dividing by $P(x)$ and taking the supremum, we obtain $\overline{N_{U^{k}}}(P)\leq\overline{N_{U}}(P)\overline{N_{U^{k-1}}}(P)$ and we conclude by recurrence. ∎
@@
First, $V^{-1}(\mathbb{R}_{+}^{*})$ is nonempty from the second statement of Def.  4. The third statement implies that $V(T^{2}(x))\leq\lambda V(T(x))$ for all $x\in\mathbb{R}^{d}$ for some $\lambda\in(0,1)$ and thus $V(T^{2}(x))>0$ implies that $V(T(x))>0$. We conclude that $V\in\mathcal{D}\left(T\right)$. The third statement also implies that $V(T(x))\leq\lambda V(x)$ for all $x\in\operatorname{dom}(V)$ for some $\lambda\in(0,1)$ and thus $N_{T}(V)\leq\lambda<1$. It can be equal to 0 if and only if for all $x$ such that $V(x)>0$, $V(T(x))=0$. ∎
Example 7 (Operator ratio for $V_{\ell}$).
We compute the operator ratio of $L$ at the Opt-Lyapunov function $V_{\ell}$ proposed in Equation (32).
$\overline{N_{L}}(V_{\ell})=\sup_{x\in{V_{\ell}}^{-1}(\mathbb{R}_{+}^{*})}\dfrac{V_{\ell}(Lx)}{V_{\ell}(x)}=\sup_{x\in V_{\ell}^{-1}(\mathbb{R}_{+}^{*})}\dfrac{x^{\intercal}x}{x^{\intercal}L^{\intercal}Lx}=\left(\inf_{x_{1}>0,\ x_{2}>0}\dfrac{x^{\intercal}L^{\intercal}Lx}{x^{\intercal}x}\right)^{-1}$
Finally, we can use a characterization of copositive matrices (see for example  [HUS10] and references therein) to compute $\overline{N_{L}}{V_{\ell}}$. Indeed, we have:
$\inf_{\begin{subarray}{c}x_{1}>0\\ x_{2}>0\end{subarray}}\dfrac{x^{\intercal}L^{\intercal}Lx}{x^{\intercal}x}=\sup\left\{\lambda\geq 0:x^{\intercal}L^{\intercal}Lx-\lambda x^{\intercal}x\geq 0\text{ on }(\mathbb{R}_{+}^{*})^{2}\right\}=\sup\left\{\lambda\geq 0:L^{\intercal}L-\lambda\mathrm{Id}\text{ is copositive}\right\}$
From [HUS10, Proposition 2.1], $L^{\intercal}Lx-\lambda\mathrm{Id}=\begin{pmatrix}a^{2}+c^{2}-\lambda&ab\\ ab&b^{2}-\lambda\end{pmatrix}$ is copositive if and only if $a^{2}+c^{2}-\lambda\geq 0$, $b^{2}-\lambda\geq 0$ and $\sqrt{(a^{2}+c^{2}-\lambda)(b^{2}-\lambda)}+ab\geq 0$. Since $a\geq 1$ and $b>1$, the third inequality is a consequence of the first two conditions. We conclude that:
$\overline{N_{L}}(V)=\sup_{x\in V^{-1}(\mathbb{R}_{+}^{*})}\dfrac{x^{\intercal}x}{x^{\intercal}L^{\intercal}Lx}=\dfrac{1}{\min\{a^{2}+c^{2},b^{2}\}}$
Applying the computations to the numerical data of Subsection 3.4, we have
$\overline{N_{L}}(V_{\ell})=\dfrac{4}{5}.$
To prove that a function $V\in\mathcal{PSD}$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, we could use any value greater $\overline{N_{T}}(V)$. The computation of $\overline{N_{T}}(V)$ ensures that the best value possible is obtained. This is important in the context of using the function $\mathfrak{F}_{\nu}(k,h,\beta)$ with an Opt-Lyapunov function because $\beta\mapsto\mathfrak{F}_{\nu}(k,h,\beta)$ is decreasing.
For classical Lyapunov functions, the extended version (on positive semidefinite functions) of operator ratios (Definition 5) coincides with the positive definite version (recalled in (17)).
@@
[…]
5.2 $(X^{\mathrm{in}},T,\varphi)$-certificate of compatibility and $(X^{\mathrm{in}},T,\varphi)$-compatibility
The construction of $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ from a classical Lyapunov function $V$ is based on the existence of a function $\alpha\in\mathcal{K}_{\infty}$ such that $\varphi\leq\alpha\circ V$. The existence of this latter function $\alpha$ is either assumed (with $\alpha=\mathrm{Id}$) in Equation (16) or in Equation (15) comes from the continuity of $\varphi$ and the definition of $V$ (see Equation (12)). However, when $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$, $V$ is neither positive definite nor continuous. Thus, we cannot guarantee the existence of $\alpha\in\mathcal{K}_{\infty}$ such that $\alpha\circ\varphi\leq V$. We must use a different class of functions to find a similar relation between $\varphi$ and $V$. First, we propose to restrict the interval in which a function $\alpha$ should be continuous and strictly increasing. Indeed, the analysis only concerns the values $\{\varphi(T^{k}(x)),k\in\mathbb{N},x\in X^{\mathrm{in}}\}$ and $\{V(x),x\in X^{\mathrm{in}}\}$. Hence, we introduce the following interval:
@@
    For all $k\in\mathbb{N}$, $\nu_{k}\in I_{X^{\mathrm{in}},T}^{\varphi}$. If the values are finite, $\limsup_{n\in\mathbb{N}}\nu_{n}$ and $\liminf_{n\in\mathbb{N}}\nu_{n}$ also belong to $I_{X^{\mathrm{in}},T}^{\varphi}$.
@@
Let $k\in\mathbb{N}$. There exists for all $n\in\mathbb{N}$, $x_{n}\in X^{\mathrm{in}}$ such that $\nu_{k}-1/n\leq\varphi(T^{k}(x_{n}))\leq\nu_{k}$ and thus the sequence $(\varphi(T^{k}(x_{n})))_{n\in\mathbb{N}}$ converges to $\nu_{k}$ and thus $\nu_{k}\in I_{X^{\mathrm{in}},T}^{\varphi}$. The values, if they are finite, $\limsup_{k\in\mathbb{N}}\nu_{k}$ and $\liminf_{k\in\mathbb{N}}\nu_{k}$ are limits of subsequences of $(\nu_{k})_{k\in\mathbb{N}}$ and they belongs to $I_{X^{\mathrm{in}},T}^{\varphi}$. From the definition of $\limsup$ (resp. $\liminf$), if $\limsup_{k\in\mathbb{N}}\nu_{k}$ (resp. $\liminf_{k\in\mathbb{N}}\nu_{k}$) is infinite, the value $\sup_{k\in\mathbb{N}}\nu_{k}$ (resp. $\inf_{k\in\mathbb{N}}\nu_{k}$ is equal to $\infty$ (resp. -$\infty$).
If one of the values $\inf_{k\in\mathbb{N}}\nu_{k}$ or $\sup_{k\in\mathbb{N}}\nu_{k}$ is not finite, the infinity passes to the related bound of $I_{X^{\mathrm{in}},T}^{\varphi}$. If those values are both finite, we can extract sequences in $I_{X^{\mathrm{in}},T}^{\varphi}$ which converge respectively to $\inf_{m\in\mathbb{N}}\nu_{m}$ and $\sup_{m\in\mathbb{N}}\nu_{m}$. For example, we can define $(\varphi(T^{k_{n}}(x_{k_{n}})))_{n}\subset I_{X^{\mathrm{in}},T}^{\varphi}$ such that for all $n\in\mathbb{N}$,
$\inf_{m\in\mathbb{N}}\nu_{m}-1/n\leq\nu_{k_{n}}-1/n\leq\varphi(T^{k_{n}}(x_{k_{n}}))\leq\nu_{k_{n}}\leq\inf_{m\in\mathbb{N}}\nu_{m}+1/n.$
This sequence converges to $\inf_{m\in\mathbb{N}}\nu_{m}$. ∎
Let $V$ be an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. To pass from $\alpha\circ\varphi\leq V$ to $\varphi\leq\displaystyle{{\alpha}_{\left(J\right)}^{-1}}\alpha\circ V$, $\alpha$ must be surjective and strictly increasing as a function from $J$ to $\{V(x),x\in X^{\mathrm{in}}\}\subseteq[0,1]$. From the earlier discussion, $J$ must also contain $I_{X^{\mathrm{in}},T}^{\varphi}$. Therefore, we define the following set of functions
@@
Let us consider $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and $I$ such that $\alpha(I)=[0,1]$. As $\alpha$ is continuous and strictly increasing on $I$, ${{\alpha}_{\left(I\right)}^{-1}}$ exists and is also continuous and strictly increasing from $[0,1]$ to $I$. Hence, $I$ is compact. Moreover, such an interval $I$ is unique. Otherwise, consider $I,J\in\mathrm{SIC}(\alpha)$ such that $I\neq J$ and $\alpha(I)=\alpha(J)=[0,1]$. Suppose there exists $x\in I$ such that $x\leq\min J$. We have $\alpha(x)<\alpha(\min J)=0$ but $x\in I$ and $\alpha(x)\geq\alpha(\min I)=0$. ∎
We present some useful properties of functions belonging to $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and their relation to $\nu$.
@@
For all $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ such that there exists an Opt-Lyapunov function $V$ for $(X^{\mathrm{in}},T)$ satisfying $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, we have $\alpha(\limsup_{k\to+\infty}\nu_{k})\leq 0$.
@@
Let $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. By assumption, $\alpha(\varphi(T^{k}(x)))\leq V(\varphi(T^{k}(x)))\leq\lambda^{k}V(x)$. Taking the supremum over $X^{\mathrm{in}}$ and from the increasing condition and the continuity of $\alpha$, $\alpha(\nu_{k})\leq\lambda^{k}\overline{V_{X^{\mathrm{in}}}}$. Taking the limsup, we get $\limsup_{k\mapsto+\infty}\alpha(\nu_{k})=\alpha(\limsup_{k\mapsto+\infty}\nu_{k})\leq 0$. ∎
@@
Suppose there exist $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ and an Opt-Lyapunov function $V$ for $(X^{\mathrm{in}},T)$ such that $\alpha(\varphi(T^{k}(x)))\leq V(T^{k}(x))$ for all $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. Then:
@@
The first statement follows readily from Prop. 10 and the fact that $\limsup_{k\to+\infty}\nu_{k}<+\infty$ iff $\nu_{\rm opt}<+\infty$. The second statement follows from Prop. 1 and Prop. 10, since we have, if ${\mathrm{R}}_{\nu}^{s}=\emptyset$, for all $x\in X^{\mathrm{in}}$ and all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\sup_{k\in\mathbb{N}}\nu_{k}=\limsup_{n\to+\infty}\nu_{n}$. ∎
The set $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$ defines the class of functions that we must consider to obtain relations $\alpha\circ\varphi\leq V$ when $V$ is an Opt-Lyapunov function for $(X^{\mathrm{in}},T)$. Another important consideration to take into account is the usefulness of functions $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ based on elements $\alpha$ of $\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$. An Opt-Lyapunov function for $(X^{\mathrm{in}},T)$ is always nonnegative and if $\alpha(\nu_{k})\leq 0$ for all $k\in\mathbb{N}$, $\alpha\circ\varphi\leq V$ obviously holds. However, a function $\alpha$ nonpositive on the terms $\nu_{k}$ whereas ${\mathrm{R}}_{\nu}^{s}$ is nonempty is not useful. To be useful, $\alpha(\nu_{k})$ must be strictly positive for some $k\in\mathbb{N}$. By Prop. 10 and Corollary 3, we have automatically $k\in{\mathrm{R}}_{\nu}^{s}$. We propose a special definition for the particular functions $\alpha\in\operatorname{SC}_{X^{\mathrm{in}},T}^{\varphi}$.